OpenStudy (rmrjr22):

Solve... y' - 2xy = e^(x^2)

OpenStudy (rmrjr22):

i ended up with ... (x+c)/(e^(-x^2)) = y

OpenStudy (tkhunny):

Well, that might be a solution if x 0 or x = -c, but with only two points of any interest, this seems unsatisfactory. How did you manage that response?

OpenStudy (tkhunny):

btw - Why would you EVER write \(\dfrac{x+c}{e^{-x^{2}}}\), when you could write \((x+c)e^{x^{2}}\)?

OpenStudy (rmrjr22):

ah that is true... thats the answer i was given... but didnt know how to get it

OpenStudy (tkhunny):

I might just be tempted to guess on thhis one. Try y = Ae^(x^2). If that result proves unsatisffaactorry, Try y = Ae^(x^2)+ Bxe^{x^2}. If yoou still don't like it,, keep adding an x and a parameter until you do like it.

OpenStudy (anonymous):

I think finding an integrating factor might require less work: \[y'-2xy=e^{x^2}\] gives you an integrating factor of \[\mu(x)=e^{\int(-2x)~dx}=e^{-x^2}\] So \[e^{-x^2}y'-2xe^{-x^2}y=e^{-x^2}e^{x^2}\\ \frac{d}{dx}\left[e^{-x^2}y\right]=1\\ e^{-x^2}y=\int dx\\ y=\frac{x+C}{e^{-x^2}}=(x+C)e^{-x^2}\]

OpenStudy (tkhunny):

Oh, okay. Fine. It's still +x^2 in the numerator.

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @SithsAndGiggles \[y=\frac{x+C}{e^{-x^2}}=(x+C)e^{x^2}\] \(\color{blue}{\text{End of Quote}}\) Yep, small typo :)